# Convex cocompact groups in real hyperbolic spaces with limit set a Pontryagin sphere

### v2

@@ 3.1. The 600-cell as a union of two solid tori
Among the six convex regular $4$-polytopes, there is precisely one with $600$ facets, namely, the hexacosichoron. From now on, we will refer to the simplicial complex given by the boundary of the hexacosichoron as the $600$-cell, denoted $\mathsf{C}_{600}$. As such, $\mathsf{C}_{600}$ is a triangulation of the 3-dimensional sphere $\mathbb{S}^{3}$ consisting of 600 tetrahedra, 1200 triangles, 720 edges, and 120 vertices. Each vertex is incident to 20 tetrahedra and 12 edges. We will now review a certain decomposition of $\mathsf{C}_{600}$ that is relevant for the purposes of this paper. For additional details about the 600-cell see [CO70, pp. 19-23, §9].
@@
    The surface $\partial T_{i}$ of each of $T_{i}$ is the same as the outer surface of the complex consisting of 10 pentagonal antiprisms—so-called drums—stacked together to form a torus with 100 exposed [^1] A simplex is exposed if it is contained in $\partial T_{i}$. triangles (with a contribution of $10$ from each drum), 150 exposed edges, and 50 exposed vertices. In particular the triangulation of $\partial T_{i}$ is that of a flat torus by 100 equilateral triangles, in which every vertex has 6 neighbors. See Figure 3.
@@
    To each of the 100 exposed triangles of $T_{0}$, we glue a raised tetrahedron, in the following way: the raised tetrahedra come in pairs, identified along the triangular faces sharing an edge that is horizontal in Figure 3. The unique common vertex among such a pair of tetrahedra which is not a vertex of $T_{0}$ will be called the raised vertex of that pair. There are 50 raised vertices. We do the same for $T_{1}$.
@@
    6 neighbors $w_{i}$ on $\partial T_{0}$. The edges $[v,w_{i}]$ are on $\partial T_{0}$.
@@
Among the six convex regular $4$-polytopes, there is precisely one with 120 facets, namely, the dodecacontachoron. The right-angled dodecacontachoron can be realized as a compact polytope $\mathcal{P}_{120}$ in $\mathbb{H}^{4}$. By Poincaré’s polyhedron theorem, the group $\mathrm{Refl}(\mathcal{P}_{120})$ generated by the reflections in the facets of $\mathcal{P}_{120}$ is a uniform lattice in $\mathbb{H}^{4}$ and $\mathrm{Refl}(\mathcal{P}_{120})$ is the RACG defined by the 1-skeleton of the dual of $\partial\mathcal{P}_{120}$, i.e., the 1-skeleton of $\mathsf{C}_{600}$, so that $\mathrm{Refl}(\mathcal{P}_{120})\cong W_{\mathsf{C}_{600}}$.
@@
The triangulation is clearly flag-no-square; indeed this is a piece of the equilateral triangulation of $\mathbb{R}^{2}$, with sides longer than $4$, see Figure 3. To see that $\partial T_{0}$ is a full subcomplex we argue as follows. Let $v$ be a vertex of $\partial T_{0}$. The above description of the neighbors of $v$ shows that all the edges of $\mathsf{C}_{600}$ connecting $v$ with another vertex on $\partial T_{0}$ are already contained in $\partial T_{0}$, namely they are the 6 edges to the 6 neighbors on $\partial T_{0}$. So, the 1-skeleton of $\partial T_{0}$ is an induced subgraph of the 1-skeleton of $\mathsf{C}_{600}$. Since $\partial T_{0}$ is a flag 2-complex, it follows that $\partial T_{0}$ is a full subcomplex. ∎
With this proposition, we can see that the limit set of $\Gamma^{4}$ is a Pontryagin sphere, as follows. Since $\partial T_{0}$ is a full subgraph of $\mathsf{C}_{600}$, it follows from (1.) in Lemma 2.1 that the special subgroup $\Gamma^{4}$ is quasiconvex and isomorphic to the abstract RACG $W_{\partial T_{0}}$. Since $\mathrm{Refl}(\mathcal{P}_{120})$ acts cocompactly on $\mathbb{H}^{4}$, Lemma 2.3 implies that $\Gamma^{4}$ is convex cocompact in $\mathrm{Isom}(\mathbb{H}^{4})$ and in particular that the limit set of $\Gamma^{4}$ is homeomorphic to the Gromov boundary of $\Gamma^{4}$. Finally, since the nerve $\partial T_{0}$ of $W_{\partial T_{0}}$ is a flag-no-square triangulation of a closed orientable surface of positive genus, the Gromov boundary of $\Gamma^{4}\cong W_{\partial T_{0}}$ is a Pontryagin sphere by Lemma 2.2.
@@
By construction, $\Gamma^{4}$ is a subgroup of a uniform lattice $\mathrm{Refl}(\mathcal{P}_{120})$ in $\mathbb{H}^{4}$. It follows from work of Bugaenko [bugaenko] that $\mathrm{Refl}(\mathcal{P}_{120})$ has finite index in the group $O(f_{\varphi},\mathbb{Z}[\varphi])$, where $\varphi=\frac{1+\sqrt{5}}{2}$, and $f_{\varphi}$ is the bilinear form on $\mathbb{R}^{5}$ given by the diagonal matrix $(1,1,1,1,-\varphi)$, so that $\mathrm{Refl}(\mathcal{P}_{120})$ is an arithmetic reflection lattice.
@@ 4.1. An abstract Coxeter group with $21$ generators
Let $T_{21}$ be the flag-no-square simplicial complex depicted in Figure 4 (vertices sharing the same label should be identified and the boundary edge should be identified accordingly), it is a simplicial complex whose geometric realization is a topological surface $S$ with 21 vertices, 63 edges and 42 faces, so that $\chi(S)=0$. Hence, $T_{21}$ is a flag-no-square triangulation of a 2-torus. The goal is to build a convex cocompact action of $W_{T_{21}}$ on $\mathbb{H}^{6}$ as a reflection group.
In fact, the presentation of $W_{T_{21}}$ can be expressed rather simply, and usefully, as follows. Let $M=(m_{i,j})$ be the symmetric $21\times 21$ matrix with entries in $\{1,2,\infty\}$ given by
@@
$A_{i,j}=\begin{cases}1&\text{if }i=j;\\ 0&\text{if }|i-j|=\pm 1,\pm 4\text{ or}\pm 5\;(\bmod\,21);\\ -u&\text{if }|i-j|=\pm 3,\pm 6\text{ or}\pm 9\;(\bmod\,21);\\ -v&\text{if }|i-j|=\pm 2,\pm 8\text{ or}\pm 10\;(\bmod\,21);\\ -w&\text{if }|i-j|=\pm 7\;(\bmod\,21),\end{cases}$
where $u=\frac{1}{50}\left(27+7\sqrt{21}\right)$, $v=\frac{1}{50}\left(21+11\sqrt{21}\right)$ and $w=\frac{1}{50}\left(49+9\sqrt{21}\right)$. A computation shows that the matrix $A$ has rank $7$ and signature $(6,1,14)$.
So, [Vin85, Theorem 2.1] shows that there is a right-angled Coxeter polytope $\mathcal{P}_{21}$ with $21$ facets in $\mathbb{H}^{6}$ whose Gram matrix is $A$. The group $\Gamma^{6}=\mathrm{Refl}(\mathcal{P}_{21})$ generated by the $21$ reflections along the boundary hyperplanes of $\mathcal{P}_{21}$ is isomorphic to $W_{T_{21}}$, by Poincaré’s polyhedron theorem.
The group $\Gamma^{6}$ is convex cocompact, since $u,v,w>1$ (see [DH13, Theorem 4.12]). Lemma 2.3.(1.) shows that the Gromov boundary of $W_{T_{21}}$ is homeomorphic to the limit set of $\Gamma^{6}$. As the nerve of $W_{T_{21}}$ is a flag-no-square triangulation of a torus, the Gromov boundary of $W_{T_{21}}$ is a Pontryagin sphere, by Lemma 2.2. Hence, the limit set of the convex cocompact subgroup $\Gamma^{6}$ is a Pontryagin sphere. ∎
@@ 4.3. A two generator group that contains $\Gamma^{6}$ as a finite-index subgroup
[…]
@@
Since the group $\Gamma^{6}$ and the group generated by $s_{1},\ldots,s_{21}$ have the same dimension and the same Gram matrix, they are conjugate by [Vin85, Theorem 2.1]. Let $\Gamma_{2}^{6}$ be the group generated by $\sigma$ and $s_{1}$. From the relations, for $k\in\{1,2,\dotsc,21\}$, $\sigma^{k}s_{1}\sigma^{-k}=s_{k}$, we get that $\Gamma^{6}$ is a normal subgroup of $\Gamma_{2}^{6}$. The index of $\Gamma^{6}$ in $\Gamma^{6}_{2}$ is equal to the order of $\sigma$ in the quotient $\Gamma_{2}^{6}/\Gamma^{6}$, hence this index is $1,3,7$ or $21$. If the order of $\sigma$ in $\Gamma_{2}^{6}/\Gamma^{6}$ is not $21$ then $\Gamma^{6}$ contains $\sigma$, $\sigma^{3}$, or $\sigma^{7}$, and hence contains an element of order $21$, $7$, or $3$. But, in a Coxeter group, a finite-order element is conjugate into a spherical special subgroup (see [DA08, Theorem. 12.3.4.(i)]). In particular, in a right-angled Coxeter group, finite-order elements are of order a power of $2$, and hence $\sigma$ must be of order $21$ in $\Gamma_{2}^{6}/\Gamma^{6}$. ∎
@@
A finite-index subgroup of a convex cocompact group is convex cocompact as well, and has the same limit set. So, Propositions 4.1 and 4.2 together imply Theorem 1.2. ∎
@@ 4.4. A few remarks
The Gram matrix of $\mathcal{P}_{21}$ contains three principal submatrices that are very symmetric. More precisely, for $j=1,2,3$, let $S_{j}=\{k\in\{1,...,21\}\,|\,k\equiv j\}$. Then the principal submatrix $A_{j}$ of the Gram matrix $A$ obtained by deleting the rows and columns not in $S_{j}$ is equal to :
@@
Geometrically, this means that the Coxeter polytope $\mathcal{P}_{21}$ is the intersection of three isometric regular simplicies $\Delta_{j}$, for $j=1,2,3$.
We claim that $\mathrm{Refl}(\Delta_{1})$ is not virtually contained in an arithmetic lattice, which trivially implies that $\mathrm{Refl}(\mathcal{P}_{21})$ is not virtually contained in an arithmetic lattice.
To that end, note first that $\mathrm{Refl}(\Delta_{1})$ is Zariski dense in $\mathrm{O}(f)$ by [benoist_harpe], where $f$ is the Tits bilinear form on $\mathbb{R}^{7}$ given by $A_{1}$. Now suppose $\mathrm{Refl}(\Delta_{1})$ were contained in an arithmetic lattice of $\mathrm{O}(f,\mathbb{R})$. Then, by [Vin71, Thm. 3], the trace of $\mathrm{Ad}(\gamma)$ would be an algebraic integer for each $\gamma\in\mathrm{Refl}(\Delta_{1})$, where $\mathrm{Ad}$ denotes the adjoint representation of $\mathrm{O}(f,\mathbb{R})$. On the other hand, since $\Gamma\subset\mathrm{O}(f,\mathbb{Q}(\sqrt{21}))$, this would imply that the trace of $\gamma$ is an algebraic integer for each $\gamma\in\mathrm{Refl}(\Delta_{1})$; see [Vin71, Theorems 1 and 2 and Lemma 2]. However, the trace of the product of two generators of $\mathrm{Refl}(\Delta_{1})$ is equal to $2(u^{2}+1)=\frac{1}{625}(3008+378\sqrt{21})$, which is not an algebraic integer. ∎
@@
Since the subgroup $\mathrm{Refl}(\Delta_{1})$ is Zariski-dense in $\mathrm{O}(f)$, we have that the group $\Gamma^{6}$ is Zariski-dense in $\mathrm{O}(f)$ as well. Hence, the limit set of $\Gamma^{6}$ is not contained in any round $k$-sphere of $\mathbb{S}^{k}=\partial\mathbb{H}^{k+1}$, for $k\leqslant 4$.

### v3

@@ 3.1. The 600-cell as a union of two solid tori
Among the six convex regular $4$-polytopes, there is precisely one with $600$ facets, namely, the hexacosichoron. From now on, we will refer to the simplicial complex given by the boundary of the hexacosichoron as the $600$-cell, denoted $\mathsf{C}_{600}$. As such, $\mathsf{C}_{600}$ is a triangulation of the 3-dimensional sphere $\mathbb{S}^{3}$ consisting of 600 tetrahedra, 1200 triangles, 720 edges, and 120 vertices. Each vertex is incident to 20 tetrahedra and 12 edges. We will now review a certain decomposition of $\mathsf{C}_{600}$ that is relevant for the purposes of this paper. For additional details about the 600-cell see [13, pp. 19-23, §9].
@@
    The surface $\partial T_{i}$ of each of $T_{i}$ is the same as the outer surface of the complex consisting of 10 pentagonal antiprisms—so-called drums—stacked together to form a torus with 100 exposed [^1] A simplex is exposed if it is contained in $\partial T_{i}$. triangles (with a contribution of $10$ from each drum), 150 exposed edges, and 50 exposed vertices. In particular the triangulation of $\partial T_{i}$ is that of a torus by 100 triangles, in which every vertex has 6 neighbors. See Figure 3.
@@
    To each of the 100 exposed triangles of $T_{0}$, we glue a tetrahedron (which we call a raised tetrahedron), in the following way: the raised tetrahedra come in pairs, identified along the triangular faces that share an edge making an angle of $60^{\circ}$ with the positive horizontal direction in Figure 3. The unique common vertex among such a pair of tetrahedra which is not a vertex of $T_{0}$ will be called the raised vertex of that pair. There are 50 raised vertices. We do the same for $T_{1}$.
@@
    6 neighbors $w_{i}$ on $\partial T_{0}$. The edges $v,w_{i}]$ are on $\partial T_{0}$.
@@
Among the six convex regular $4$-polytopes, there is precisely one with 120 facets, namely, the dodecacontachoron. The right-angled dodecacontachoron can be realized as a compact polytope $\mathcal{P}_{120}$ in $\mathbb{H}^{4}$ (see e.g. [12, 18] or [39, Section 1.3]). By Poincaré’s polyhedron theorem, the group $\mathrm{Refl}(\mathcal{P}_{120})$ generated by the reflections in the facets of $\mathcal{P}_{120}$ is a uniform lattice in $\mathbb{H}^{4}$ and $\mathrm{Refl}(\mathcal{P}_{120})$ is the RACG defined by the 1-skeleton of the dual of $\partial\mathcal{P}_{120}$, i.e., the 1-skeleton of $\mathsf{C}_{600}$, so that $\mathrm{Refl}(\mathcal{P}_{120})\cong W_{\mathsf{C}_{600}}$.
@@
The triangulation is clearly flag-no-square; indeed this is a quotient of an equilateral triangulation of a parallelogram in $\mathbb{R}^{2}$, whose sides have combinatorial length larger than $4$, see Figure 3. To see that $\partial T_{0}$ is a full subcomplex we argue as follows. Let $v$ be a vertex of $\partial T_{0}$. The above description of the neighbors of $v$ shows that all the edges of $\mathsf{C}_{600}$ connecting $v$ with another vertex on $\partial T_{0}$ are already contained in $\partial T_{0}$, namely they are the 6 edges to the 6 neighbors on $\partial T_{0}$. So, the 1-skeleton of $\partial T_{0}$ is an induced subgraph of the 1-skeleton of $\mathsf{C}_{600}$. Since $\partial T_{0}$ is a flag 2-complex, it follows that $\partial T_{0}$ is a full subcomplex. ∎
With this proposition, we can see that the limit set of $\Gamma^{4}$ is a Pontryagin sphere, as follows. Since $\partial T_{0}$ is a full subcomplex of $\mathsf{C}_{600}$, it follows from (1.) in Lemma 2.1 that the special subgroup $\Gamma^{4}$ is quasiconvex and isomorphic to the abstract RACG $W_{\partial T_{0}}$. Since $\mathrm{Refl}(\mathcal{P}_{120})$ acts cocompactly on $\mathbb{H}^{4}$, Lemma 2.3 implies that $\Gamma^{4}$ is convex cocompact in $\mathrm{Isom}(\mathbb{H}^{4})$ and in particular that the limit set of $\Gamma^{4}$ is homeomorphic to the Gromov boundary of $\Gamma^{4}$. Finally, since the nerve $\partial T_{0}$ of $W_{\partial T_{0}}$ is a flag-no-square triangulation of a closed orientable surface of positive genus, the Gromov boundary of $\Gamma^{4}\cong W_{\partial T_{0}}$ is a Pontryagin sphere by Lemma 2.2.
@@
By construction, $\Gamma^{4}$ is a subgroup of a uniform lattice $\mathrm{Refl}(\mathcal{P}_{120})$ in $\mathbb{H}^{4}$. It follows from work of Bugaenko [9] that $\mathrm{Refl}(\mathcal{P}_{120})$ has finite index in the group $O(f_{\varphi},\mathbb{Z}[\varphi])$, where $\varphi=\frac{1+\sqrt{5}}{2}$, and $f_{\varphi}$ is the bilinear form on $\mathbb{R}^{5}$ given by the diagonal matrix $(1,1,1,1,-\varphi)$, so that $\mathrm{Refl}(\mathcal{P}_{120})$ is an arithmetic reflection lattice.
@@ 4.1. An abstract Coxeter group with $21$ generators
Let $T_{21}$ be the flag simplicial complex determined by the graph depicted in Figure 4 (vertices sharing the same label should be identified and the boundary edges should be identified accordingly). Since every homotopically nontrivial edge-cycle in $T_{21}$ has combinatorial length greater than $4$, the complex is flag-no-square. The geometric realization of $T_{21}$ is a topological surface $S$ with 21 vertices, 63 edges and 42 faces, so that $\chi(S)=0$. Hence, $T_{21}$ is a flag-no-square triangulation of a 2-torus. The goal is to build a convex cocompact action of $W_{T_{21}}$ on $\mathbb{H}^{6}$ as a reflection group.
In fact, the presentation of $W_{T_{21}}$ can be expressed rather simply, and usefully, as follows. For $i=1$, the vertices adjacent to the vertex labeled $i$ in $T_{21}$ have the labels $i\pm 1$, $i\pm 4$, and $i\pm 5$ (mod $21$). Additionally, when one goes from the vertex $i$ to the vertex $i+1$, the labels of all adjacent vertices increase by $1$ (mod $21$). Hence, let $M=(m_{i,j})$ be the symmetric $21\times 21$ matrix with entries in $\{1,2,\infty\}$ given by
@@
[…]
where $u=\frac{1}{50}\left(27+7\sqrt{21}\right)$, $v=\frac{1}{50}\left(21+11\sqrt{21}\right)$ and $w=\frac{1}{50}\left(49+9\sqrt{21}\right)$. A computation shows that the matrix $A$ has rank $7$ and signature $(6,1,14)$, see Remark 4.2. So, [50, Theorem 2.1] shows that there exists a right-angled Coxeter polytope $\mathcal{P}_{21}$ with $21$ facets in $\mathbb{H}^{6}$ whose Gram matrix is $A$. The group $\Gamma^{6}=\mathrm{Refl}(\mathcal{P}_{21})$ generated by the $21$ reflections along the boundary hyperplanes of $\mathcal{P}_{21}$ is isomorphic to $W_{T_{21}}$, by Poincaré’s polyhedron theorem, since $u,v,w\geqslant 1$.
The group $\Gamma^{6}$ is convex cocompact since $\mathcal{P}_{21}$ has no asymptotic facets, as $u,v,w>1$ (see [20, Theorem 4.12]). Lemma 2.3.(1.) shows that the Gromov boundary of $W_{T_{21}}$ is homeomorphic to the limit set of $\Gamma^{6}$. As the nerve of $W_{T_{21}}$ is a flag-no-square triangulation of a torus, the Gromov boundary of $W_{T_{21}}$ is a Pontryagin sphere, by Lemma 2.2. Hence, the limit set of the convex cocompact subgroup $\Gamma^{6}$ is a Pontryagin sphere. ∎
Remark 4.2.
The above values for $u,v,w$ were obtained as follows. We initialized a $21\times 21$ matrix $A$ with variable entries $u,v,w$, and looked for values for $u,v,w$ at which $A$ has rank $7$ and signature $(6,1,14)$. To this end, we sampled some $8\times 8$ minors and imposed that their determinant be zero. This provided a system of polynomial equations in $u,v,w$. First, we solved these numerically, then the LLL method was used to obtain algebraic numbers of bounded degree close to the numerical solutions. These are the values for $u,v,w$ written above, and we verified a posteriori that the matrix obtained for those specific values had the desired signature. This was done by an exact algebraic computation with SageMath. Detailed computations are provided in the SageMath file (computations.ipynb), available on the webpage [21]. The fact that only three variables sufficed is of course surprising, but probably is a consequence of the fact that $T_{21}$ is very symmetric.
@@ 4.3. A two generator group that contains $\Gamma^{6}$ as a finite-index subgroup
Let $y=\left(\alpha,0,\alpha,0,\alpha,0,\beta\right)$, where $\alpha=\tfrac{1}{5}\sqrt{11+\sqrt{21}}$ and $\beta=\tfrac{1}{5}\sqrt{8+3\sqrt{21}}$. Note that $y$ is a space-like vector of $\mathbb{R}^{6,1}$ with length $1$, that is, $\langle y,y\rangle=1$, where $\langle\phantom{y},\phantom{y}\rangle$ is the standard Minkowski metric on $\mathbb{R}^{6,1}$. Therefore, for each integer $k\in\{1,2,\dotsc,21\}$, a reflection $s_{k}\in\mathrm{O}(6,1)$ can be defined by
@@
Again, detailed computations can be found in the SageMath file (computations.ipynb), available on the webpage [21].
Since the group $\Gamma^{6}$ and the group generated by $s_{1},\ldots,s_{21}$ both lie in $\mathrm{O}(6,1)$ and share the same Gram matrix, they are conjugate by [50, Theorem 2.1]. Let $\Gamma_{2}^{6}$ be the group generated by $\sigma$ and $s_{1}$. From the relations, for $k\in\{1,2,\dotsc,21\}$, $\sigma^{k-1}s_{1}\sigma^{-k+1}=s_{k}$, we get that $\Gamma^{6}$ is a normal subgroup of $\Gamma_{2}^{6}$. The index of $\Gamma^{6}$ in $\Gamma^{6}_{2}$ is equal to the order of $\sigma$ in the quotient $\Gamma_{2}^{6}/\Gamma^{6}$, hence this index is $1,3,7$ or $21$. If the order of $\sigma$ in $\Gamma_{2}^{6}/\Gamma^{6}$ is not $21$ then $\Gamma^{6}$ contains $\sigma$, $\sigma^{3}$, or $\sigma^{7}$, and hence contains an element of order $21$, 77, or $3$. But, in a Coxeter group, a finite-order element is conjugate into a spherical special subgroup (see [19, Theorem. 12.3.4.(i)]). In particular, in a right-angled Coxeter group, finite-order elements are of order a power of $2$, and hence $\sigma$ must be of order $21$ in $\Gamma_{2}^{6}/\Gamma^{6}$. ∎
@@
A finite-index subgroup of a convex cocompact group is convex cocompact as well, and has the same limit set. So, Propositions 4.1 and 4.3 together imply Theorem 1.2. ∎
@@ 4.4. A few remarks
The Gram matrix of $\mathcal{P}_{21}$ contains three principal submatrices that are very symmetric. More precisely, for $j=1,2,3$, let $S_{j}=\{k\in\{1,...,21\}\,|\,k\equiv j\,\,(\mathrm{mod}\,3)\}$. Then the principal submatrix $A_{j}$ of the Gram matrix $A$ obtained by deleting the rows and columns not in $S_{j}$ is equal to :
@@
Each $A_{j}$ has signature $(6,1,0)$. Geometrically, the Coxeter polytope $\mathcal{P}_{21}$ is the intersection of three isometric polyhedra $\Delta_{j}$ of $\mathbb{H}^{6}$, for $j=1,2,3$, corresponding to the submatrices $A_{j}$. Each $\Delta_{j}$ is the intersection with $\mathbb{H}^{6}$ (in the projective model) of a simplex in $\mathbb{RP}^{6}$ all of whose facets have nonempty intersection with $\mathbb{H}^{6}$.
Let $f$ be the bilinear form on $\mathbb{R}^{7}$ defined by the symmetric matrix $A_{1}$ (sometimes called the Tits bilinear form associated with $A_{1}$ in the context of reflection groups), and let $\mathrm{Refl}(\Delta_{1})$ be the subgroup of $\mathrm{O}(f,\mathbb{R})$ generated by reflections across the hyperplanes orthogonal to the canonical basis vectors of $\mathbb{R}^{7}$ with respect to the bilinear form $f$. We claim that $\mathrm{Refl}(\Delta_{1})$ is not virtually contained in an arithmetic lattice of $\mathrm{O}(f,\mathbb{R})$, which implies that $\mathrm{Refl}(\mathcal{P}_{21})$ is not virtually contained in an arithmetic lattice.
To that end, note that any $\mathrm{Refl}(\Delta_{1})$-invariant subspace of $\mathbb{R}^{7}$ either is contained in the intersection of the reflection hyperplanes or contains the subspace spanned by their dual vectors (see e.g. [49, Proposition 19] or [14, Proposition 3.23]). Since $A_{1}$ has full rank, any such subspace is necessarily trivial. Hence, the action of $\mathrm{Refl}(\Delta_{1})$ on $\mathbb{R}^{7}$ is irreducible. Since $f$ is Lorentzian, any irreducible subgroup of $\mathrm{O}(f,\mathbb{R})$ has Zariski closure containing $\mathrm{SO}(f,\mathbb{R})$ (see e.g. [5, Proposition 1]). Therefore, $\mathrm{Refl}(\Delta_{1})$ is Zariski-dense in $\mathrm{O}(f,\mathbb{R})$.
Now suppose for a contradiction that there exists a finite index subgroup $\Gamma_{1}$ of $\mathrm{Refl}(\Delta_{1})$ that is contained in an arithmetic lattice $\Lambda$ of $\mathrm{O}(f,\mathbb{R})$. It is classical that, in the language of [49], one obtains from the definition of an arithmetic lattice (see e.g. [38, §IX.1] and [49, Theorem 2]) that the group $\mathrm{Ad}(\Lambda)$ is definable over $\mathcal{A}$, where $\mathcal{A}$ denotes the ring of real [^2] The fact that $\mathscr{A}$ can be taken to be the ring of real algebraic integers, as opposed to the entire ring of algebraic integers, uses that $\mathrm{O}(f,\mathbb{R})$ is not locally isomorphic to a complex Lie group. In fact, the arithmetic origin of all arithmetic lattices in $\mathrm{O}(2k,1)$ for $k\geqslant 1$, as well as all arithmetic lattices in $\mathrm{O}(n,1)$ containing Zariski-dense subgroups of $\mathrm{SO}(n-1,1)$ for any $n\geqslant 2$, is completely straightforward; such lattices are all of simplest type (see [3] and the references therein). algebraic integers and $\mathrm{Ad}$ denotes the adjoint representation of $\mathrm{O}(f,\mathbb{R})$. So $\mathrm{Ad}(\Gamma_{1})$ is definable over $\mathcal{A}$ as well. It then follows from [49, Theorem 3] that $\mathrm{Ad}(\mathrm{Refl}(\Delta_{1}))$ is also definable over $\mathcal{A}$. On the other hand, since each standard generator $s:\mathbb{R}^{7}\rightarrow\mathbb{R}^{7}$ of $\mathrm{Refl}(\Delta_{1})$ is given, using the standard basis vector $e_{s}\in\mathbb{R}^{7}$, by $v\mapsto v-(v^{T}A_{1}e_{s})$, we have that $\mathrm{Refl}(\Delta_{1})\subset\mathrm{O}(f,\mathbb{Q}(\sqrt{21}))\subset\mathrm{O}(f,\overline{\mathbb{Q}}\cap\mathbb{R})$, and hence we obtain again from [49, Theorem 2] that $\mathrm{Refl}(\Delta_{1})$ is itself definable over $\mathcal{A}$. In particular, by [49, Lem. 1], we have that the trace of each element of $\mathrm{Refl}(\Delta_{1})$ is an algebraic integer. However, the trace of the product of two generators of $\mathrm{Refl}(\Delta_{1})$ is equal to $4u^{2}+3=\frac{21}{625}(173+18\sqrt{21})$, which is not an algebraic integer. ∎
@@
Since the subgroup $\mathrm{Refl}(\Delta_{1})$ is Zariski-dense in $\mathrm{O}(f,\mathbb{R})$, we have that the group $\Gamma^{6}$ is Zariski-dense in $\mathrm{O}(f,\mathbb{R})$ as well. Hence, the limit set of $\Gamma^{6}$ is not contained in any round $k$-sphere of $\mathbb{S}^{k}=\partial\mathbb{H}^{k+1}$, for $k\leqslant 4$.